# Approaching the thermodynamic limit of a bounded one-component plasma

### v1

@@ V Conclusion
In this study, we employed the model of BOCP to investigate the thermodynamic limit of a system of ions on the compensating background without introducing the PBC. This objective was reached due to discovered universal form of the size dependence of system energy, which proved to be accurate enough to be extrapolated to infinite number of the ions in a wide range of the Coulomb coupling parameter from 0.03 to 1000. This enabled us to avoid the Ewald summation procedure and estimate the energies of a system, which is free from the additional strong interionic correlations imposed by the PBC. A very small width of the metastable region near the BOCP melting point makes it possible to evaluate its position, which amounts to $\Gamma_{\mathrm{m}}=174\pm 2$. Also, we have introduced three converging quantities in addition to the system energy, namely, the excess interatomic electrostatic energy, the excess ion–background electrostatic energy, and the ionic pressure (compressibility factor), which we determined in the MD simulation. These quantities satisfy the limit conditions that follow from the theory. We show that neither the ionic pressure can be obtained from the energy of entire system nor the excess pressure of the entire system can be estimated based on the virial determined from MD or MC simulations. We propose an accurate approximation of the ionic compressibility factor and compare it with that calculated for OCP using LAMMPS. We have established that at certain cutoff parameter $r_{c}$ introduced in LAMMPS, which is characteristic of the ionic interactions, the virial for OCP coincides with that for BOCP. We argue that this is a true virial ensured by a correct choice of $r_{c}$, which turns out to be a function of $\Gamma$. We present a wide-range approximation of this function for applications. Since $r_{c}$ defines interionic forces, its correct choice is of crucial importance for the interfacial and dynamic properties of the Coulomb systems including the transport ones. As an example, we simulate the metastable region in the neighborhood of the system melting point with a true $r_{c}$ and that from the LAMMPS manual. The results are indicative of the fact that for a true $r_{c}$, the width of metastable region increases considerably, which points to the increase in the interfacial surface tension. We stress that in every OCP simulation, the obtained virial must be in compliance with the ionic compressibility factor listed in Table 2 and shown in Fig. 9. Judging from Fig. 10, such compliance must be most important at high $\Gamma$.
@@
Apart from revision of our BOCP data for high $\Gamma$, an urgent problem to be addressed in the future is development of a calculation technique that would make it possible to determine the excess interatomic and ion–background energy directly in the simulation of OCP with PBC. Also, calculation of the ionic pressure tensor and phonon spectrum based on these energies are of undoubtful interest.
@@ Acknowledgments
This work was supported by the Ministry of Science and Higher Education of the Russian Federation (State Assignment No. 075-00270-24-00).

### v2

@@ VI Conclusion
In this study, we employed the model of BOCP to investigate the thermodynamic limit of a system of ions on the compensating background without introducing the PBC. This objective was reached due to the discovered universal form of the size dependence of system energy, which proved to be accurate enough to be extrapolated to an infinite number of the ions in a wide range of the Coulomb coupling parameter from 0.03 to 1000. This enabled us to avoid the Ewald summation procedure and estimate the characteristic energies of the system, which are free from additional strong interionic correlations imposed by the PBC. A very small width of the metastable region near the BOCP melting point makes it possible to evaluate its position, which is $\Gamma_{\mathrm{m}}=174\pm 2$. Also, we have introduced three converging quantities in addition to the system energy, namely, the excess interatomic electrostatic energy, the excess ion–background electrostatic energy, and the ionic pressure (compressibility factor), which we determined in the MD simulation. These quantities satisfy the limit conditions that follow from the theory. We show that neither the ionic pressure can be obtained from the energy of the entire system nor the excess pressure of the entire system can be estimated based on the virial determined from MD or MC simulations. We propose an accurate approximation of the ionic compressibility factor and compare it with that calculated for OCP using LAMMPS. We have established that at a certain value of the cutoff parameter $r_{c}$ introduced in LAMMPS, which is characteristic of the effective ionic interactions, the virial for OCP coincides with that for BOCP. We argue that this is a true virial ensured by a correct choice of $r_{c}$, which turns out to be a function of $\Gamma$. We present a wide-range approximation of this function for applications. Our data demonstrate that the system electrostatic energy is almost insensitive to $r_{c}$, while the ionic compressibility factor exhibits a strong dependence on this parameter. This indicates that convergence-acceleration schemes reproduce the energy landscape well but are much less accurate for the forces. Thus, the force-based quantities for the classical Coulomb OCP may become ambiguous.
Since $r_{c}$ defines the effective interionic forces, its correct choice is of crucial importance for the interfacial and dynamic properties of the Coulomb systems. As an example, we simulate the metastable region in the neighborhood of the system melting point with a true $r_{c}$ and the most commonly used. The results are indicative of the fact that for a true $r_{c}$, the width of the metastable region increases considerably, which points to the increase in the interfacial surface tension. We stress that in a correct classical OCP simulation, the obtained virial must be in compliance with the ionic compressibility factor listed in Table 2 and shown in Fig. 9. Judging from Fig. 10, such compliance must be most important at high $\Gamma$.
@@
Apart from a revision of our BOCP data for high $\Gamma$, an urgent problem to be addressed in the future is development of a calculation technique that would make it possible to determine the excess interatomic and ion–background energy directly in the simulation of OCP with PBC. Also, calculations of the ionic pressure tensor, phonon spectrum, and transport properties based on these energies are of undoubted interest.
@@ Acknowledgments
This work was supported by the Ministry of Science and Higher Education of the Russian Federation (State Assignment No. 075-00270-26-$0$).